OpenStudy (anonymous):

x^3/5 over x^1/4 x^3/5 over x^1/4 @Mathematics

OpenStudy (anonymous):

x^(7/20)

OpenStudy (goldrush18):

\[x^3/^5 \div x^1/^4=x^3/^5-^1/^4=x^7/20\]